# A Large-scale Empirical Study on Large Language Models for Election Prediction

### v1

@@ 2.2.3 Version 3: Multi-step Reasoning with Contextual Information
Applying the multi-step method at the state level for the 2020 election, we evaluated its performance across a range of scenarios, including “deep red” states (e.g., Alabama), “deep blue” states (e.g., California), and 11 swing and tipping-point states (details in Appendix C). This method demonstrated significantly improved accuracy, with predicted vote distributions closely aligning with actual results across almost all Chosen states.
Given these accuracy gains and enhanced alignment with real-world data, we adopt Version 3 as our final pipeline for election prediction. By structuring the reasoning process into multiple steps, this approach effectively mitigates skewness and captures the complex dynamics of voter behavior across diverse state contexts.
2.3 Empirical Validation and Comparative Analysis of the Proposed Pipelines
@@ 2.3.1 Public ANES Benchmark Evaluation
We first evaluated the pipelines using the ANES 2016 and 2020 Time Series datasets Studies (2019, 2022) to assess their overall performance and to verify V3’s ability to generate the Conservative-Liberal Spectrum feature. For V1 (Demographic-only Prompting) and V2 (Single-step Prompting with Time-based Information), we excluded the ideological spectrum, relying solely on demographic data. In contrast, V3 employs multi-step reasoning, inspired by Chain of Thought prompting Wei et al. (2022), to incorporate ideological alignment alongside demographics and time-sensitive factors. While the 2020 dataset lacked certain demographic variables—requiring restoration of the spectrum—the 2016 dataset allowed a direct comparison between LLM-generated (3rd Pipeline_G) and restored (3rd Pipeline_featureBack) spectra.
@@
Figure 3: Comparison of the three pipelines on ANES 2016 and 2020. The y-axis shows the predicted ratio (Eq. 1), with 0.5 indicating balance. The red baseline represents ground truth ratios from the ANES dataset.
As shown in Fig. 3, excluding the spectrum-skewed predictions toward the election winner, V1 and V2 favored Republicans in 2016 (63.25%, 66.38% support for Trump) and Democrats in 2020 (30.14%, 10.02%). By contrast, V3 closely matched the actual results (e.g., 48.38% vs. an actual 47.7% for Trump in 2016). Notably, the LLM-generated spectrum in V3 slightly outperformed the restored version, showing the model’s ability to produce meaningful ideological features. These highlight the importance of incorporating ideological alignment to improve accuracy and reduce bias.
2.3.2 2020 State-Level Evaluation
State-level simulations using synthetic data were conducted to compare the model predictions against the official 2020 election outcomes reported by the Federal Election Commission (FEC). We selected five red states, five blue states, and 11 swing or tipping-point states. The V1 pipeline (Demographic-only Prompt) displayed a consistent bias toward the Democratic Party, even in historically Republican states such as South Carolina (SC) and Ohio (OH), underscoring the limitations of relying solely on demographic features. Introducing time-dependent, election-year-specific information in V2 (Time-dependent Prompt) reduced this bias, but the predictions still skewed Democratic.
In contrast, the V3 pipeline (Multi-step Reasoning) demonstrated the highest accuracy and effectively mitigated bias, closely aligning its predictions with real outcomes. It correctly forecasted the results in 9 out of 11 swing states, with only minor misestimations in North Carolina (NC) and Arizona (AZ), and accurately captured voter behavior in deeply red and blue states. These findings highlight V3’s capability to model complex voter dynamics and closely track actual election trends. Detailed results are provided in Appx. C.
2.3.3 2024 State-Level Evaluation
Before the official 2024 results were available, we employed V3 to forecast outcomes across all 50 states. For consistency and cost-efficiency, V1 and V2 were evaluated on the same subset of states. This analysis tested V3’s predictive performance in an unseen future context and allowed for comparisons with previous research Zhang et al. (2024a).
V3 performed robustly in 2024, which accurately predicted outcomes in all traditional red and blue states (with the exception of Alaska) and successfully forecasted Trump’s wins in Arizona (AZ), Wisconsin (WI), Florida (FL), and Texas (TX), as well as Harris’s victories in New Hampshire (NH) and Minnesota (MN). Although V3 slightly mispredicted several swing states, these errors were significantly smaller than the pronounced Democratic bias exhibited by V1 and V2.
Overall, these evaluations confirm V3’s strong ability to simulate voter behavior, reduce political bias, and improve alignment with real-world election patterns. Further details on the 2024 predictions are presented in Appx. D.
3 Beyond Accuracy: Insights and Future Directions Beneath Results
In §2, we proposed a voting prediction pipeline (§2.2.3) that demonstrated high accuracy using the ratio metric Eq. (1). This section delves deeper by introducing additional evaluation metrics and analyzing the broader implications of the results.
Using these metrics, we revisit the 2020 and 2024 state-level predictions to evaluate pipelines’ overall accuracy and examine systematical biases toward political parties (§3.1). Finally, we analyze how LLM simulations align with real-world demographic trends and discuss strategies for mitigating bias and enhancing the robustness of future LLM-based election forecasting methods (§3.2).
Additional Evaluation Metrics. We consider additional three evaluation metrics to assess our predictions against actual election results: Weighted Absolute Error (WAE) (Eq. (A4)), Weighted Mean Squared Error (WMSE) (Eq. (A5)), and Bias Metric (BM) (Eq. (A6)). These metrics provide a comprehensive evaluation of model performance. WAE and WMSE measure the degree of alignment between predicted proportions and actual outcomes, with WMSE placing greater emphasis on larger errors due to its squared formulation. BM quantifies systematic biases, indicating whether the predictions consistently favor one party over the other. Together, these metrics offer a nuanced understanding of both the magnitude of prediction errors and the potential biases within the pipelines. See Appx. B.2 for details.
3.1 Result Analysis on Systematical Accuracy and Bias
To align our evaluation of the 2020 and 2024 pipelines with other studies, we selected 11 swing and tipping-point states as the scope for analysis.
Metric Vanilla (%) V2 (%) V3 (%)
Weighted Absolute Error 22.78 14.97 5.24
Weighted Mean Squared Error 5.46 2.34 0.37
@@ 3.1 Result Analysis on Systematical Accuracy and Bias
Table 1: Evaluation metrics for the 2020 election simulation across three pipelines (Vanilla, V2, and V3). V3 achieves the lowest errors and minimal bias, showing substantial improvement over the simpler approaches.
Metric Vanilla (%) V2 (%) V3 (%) C (%)
Weighted Absolute Error 21.35 25.96 3.49 4.70
Weighted Mean Squared Error 4.83 6.89 0.22 0.30
Bias Metric (BM) -21.35 -25.96 -2.95 1.26
Table 2: Evaluation metrics for the 2024 election simulation comparing three pipeline versions (Vanilla, V2, V3) and one external comparison model (C, from Zhang et al. (2024a)). As with 2020, V3 achieves the most accurate results (lowest WAE and WMSE). Although BM remains slightly negative for V3, it is greatly reduced compared to the Vanilla and V2 pipelines, indicating a less biased prediction. The comparison model (C) performs competitively, but does not surpass V3.
Results Analysis of 2020. We begin our extended analysis by examining the model performances on the 2020 election. Table 1 presents the results for three pipelines: Vanilla (V1), V2, and V3. Here, WAE and WMSE serve as measures of how closely predicted vote proportions match the actual results, while the BM indicates the presence and direction of any systematic lean. For the 2020 simulation, the Vanilla pipeline shows substantial bias towards Democratic Party, as evidenced by its large negative BM value and correspondingly high error rates. Incorporating time-dependent information in V2 reduces both WAE and WMSE compared to Vanilla, but still a notable Democratic skew persists, indicated by a still-negative BM. In contrast, V3 significantly lowers both absolute and squared errors and nearly eliminates bias, achieving a BM close to zero. This improvement confirms that the multi-step reasoning approach employed by V3—incorporating ideological placement and time-sensitive data—effectively counters the model’s initial skew and enhances overall accuracy.
Results Analysis of 2024. We next turn to the 2024 predictions. Table 2 compares the same three pipelines with an additional external comparison model (C) from Zhang et al. (2024a). The pattern observed in 2020 holds true in 2024: Vanilla and V2 again exhibit marked bias toward the Democratic Party, reflected in large negative BM values. Although V3 still shows a slight Democratic lean (BM = -2.95), it is significantly less than in the simpler pipelines. V3 also attains the lowest WAE and WMSE, outperforming both the earlier pipelines and the comparison model C. Not only does this suggest that the design principles in V3 generalize well across different election years, but it also demonstrates that structured reasoning, richer contextual inputs, and ideological self-placement can improve forecasting accuracy and reduce skew even when compared to other leading approaches.
@@
Future Direction 1: Addressing Embedded Political Skewness in Pretrained Corpora. The persistent Democratic skew in simpler pipelines and the residual bias in V3 hint at deeper sources of imbalance in the pretrained corpus. These biases could arise from uneven coverage of political perspectives or disproportionate representation of certain ideologies in the training data Jenny et al. (2024). Addressing this challenge involves examining the corpus composition, adopting balanced data selection strategies, and exploring model-level techniques—such as adversarial debiasing or targeted prompt engineering—to mitigate inherent skew. By addressing these root causes, future election forecasting models can achieve more balanced and trustworthy predictions, ultimately providing a more reliable tool for political analysis and decision-making Li et al. (2024a).
3.2 Result Analysis on LLM-Simulated Voting Patterns
Beyond examining aggregated errors and directional biases, it is important to determine whether the LLM’s predictions capture the demographic patterns observed in real-world voting behavior. We focus on four demographic dimensions—gender, ethnicity, age, and education—identified by Pew Research Center’s 2020 study, Behind Biden’s 2020 Victory Center (2021), as pivotal factors influencing voter preferences. To facilitate this comparison, we aggregated over 200,000 simulated personas across all selected states in 2020, categorizing each persona as either Republican or Democratic based on their predicted votes (excluding non-preference votes). We then examined how these personas’ preferences aligned with the demographic patterns reported by Pew.
(a) Gender and Race Gap: LLM Simulated vs. Pew Report
(b) The Age Gap: LLM Simulated vs. Pew Report
(c) The Education Gap: LLM Simulated vs. Pew Report
Figure 4: LLM-simulated voting gaps across various demographic dimensions: race and gender (a), age (b), and education (c). Each subfigure shows how the LLM’s predictions vary across key demographic factors.
Figure 4 illustrates these comparisons. Directionally, the LLM aligns with Pew’s observations: men, white voters, older adults, and those with lower educational attainment lean more Republican, mirroring established voting patterns. However, the magnitude of these differences is substantially overstated. For instance, while Pew’s data suggest that males only had about a 4% margin in favor of Trump and females about an 11% margin in favor of Biden, the LLM’s simulations inflate these margins to 47.6% among males and 45.6% among females. Similarly, Pew’s data show more modest differences across racial and educational groups, whereas the LLM produces far larger gaps, implying a tendency to reinforce stereotypes rather than reflect real-world proportions.
@@ 3.2 Result Analysis on LLM-Simulated Voting Patterns
Future Direction 2: Mitigating the Exacerbation of Stereotypical Biases. While the LLM’s ability to replicate the directional trends found in real-world data is encouraging, the large overemphasis on these demographic distinctions raises concernsChang et al. (2024). Such exaggerations risk reinforcing harmful stereotypes and potentially misrepresenting demographic groups. Moving forward, research should concentrate on developing techniques to calibrate LLM outputs, refine prompt designs, or incorporate counterbalancing information, all aimed at producing simulations that are both directionally accurate and proportionally realisticPark et al. (2024). Ensuring that LLM-based electoral forecasts maintain fairness, rather than amplifying stereotypes, is a critical step toward building more ethical and reliable computational social science tools.
@@ 4 Related Work
Applying LLMs in political science represents a new and rapidly evolving field. Although LLMs have revolutionized natural language processing and a variety of other domains, their potential in political science remains largely untapped. Initial studies have demonstrated promising results in election forecasting, policy analysis, and public opinion simulation Smith and Doe (2023); Johnson and Lee (2024). However, political science involves complex social dynamics and multilayered causal relationships, which pose significant challenges to effectively utilize LLMs Brown (2023). Recent work, e.g., Chen and Rodriguez (2024), emphasizes the need for domain-specific fine-tuning to accurately capture the nuances of political discourse, highlighting the limitations of general-purpose LLMs.
LLMs’ applications in election forecasting is still in its early stages. For instance, the “Political Campus” project by Roberts et al. (2023) created a benchmark dataset for evaluating LLM performance on election-related tasks, highlighting both potential and limitations in political forecasting. Similarly, the work by Kim and Patel (2024) using LLMs for policy sentiment analysis underscores the need for careful interpretation of results due to the challenges associated with the inherent complexity of political language. Our research aims to address these challenges by integrating LLMs with ABMs to predict election outcomes and comprehensively analyze voter behavior. This integrated framework combines macro-level political discourse analysis with micro-level voter behavior modeling, bridging the gap identified by previous studies such as Lopez and Singh (2023) regarding the disconnect between large-scale text analysis and individual-level decision-making. For a more comprehensive related work, please see Appx. A.
5 Conclusion and Future Directions
In this work, we present a novel framework for election prediction using large language models (LLMs) with a focus on multi-step reasoning. By leveraging both synthetic personas and real-world datasets, we demonstrated the potential of LLMs to capture individual voting behaviors and state-level election outcomes. Our iterative design highlights the importance of integrating temporal information and complex reasoning for accurate predictions. The 2020 and 2024 simulations reveal both the strengths and limitations of using LLMs in dynamic political environments. These results highlight the model’s ability to generalize on unseen data, while also underscoring several challenges, including: (1) evaluating and addressing biases embedded within the LLM’s pretrained corpus, (2) mitigating stereotypical and overly simplistic representations in persona simulations, and (3) overcoming static demographic assumptions and the lack of real-time data integration.
Future research can extend this work by incorporating multiple LLMs to better understand their internal political tendencies, enhancing temporal modeling with public opinion data and real-time trends for improved accuracy, and developing stronger multi-step reasoning pipelines through refined Chain of Thought (CoT) designs to further enhance prediction performance and mitigate biases. This study lays the foundation for future applications of LLMs in political forecasting, offering promising directions for further development in both election prediction and the broader study of LLM behavior in social contexts.
Broader Impact Statement
This work explores the application of LLMs to improve election forecasting through enhanced reasoning and data synthesis, which can inform policymakers, researchers, and journalists, aiding them in understanding voter behavior and electoral outcomes. By offering a more transparent and adaptable approach to prediction, this research may help demystify complex political processes, reduce reliance on narrow historical data, and guide strategic resource allocation for stakeholders.

### v3

@@ 2.2.3 V3: Multi-Step Reasoning with Ideology Inference
Our theory-driven multi-step pipeline significantly improves the LLM’s ability to model real voting behavior in both public benchmark validation and state-level simulations. Therefore, we adopt V3 as our final pipeline for voter behavior simulation. By structuring reasoning into multiple steps, this approach helps mitigate bias and better captures voter dynamics across diverse states. In the following section, we will provide a detailed discussion of our simulation results.
2.3 Empirical Validation and Cross-Model Evaluation of the Proposed Pipelines
@@ 2.3.1 Public ANES Benchmark Evaluation
We first validate our proposed framework using GPT-4o on ANES 2016 and 2020 Time Series datasets Studies (2019, 2022), which include demographic information, political ideology, and actual voting records from human respondents. Testing our models on these public benchmarks allows for a direct comparison between LLM-generated predictions and real-world human voting behavior.
As shown in Fig. 2, we assess our three pipeline versions on ANES. V1 (§2.2.1): Demographic-Only Prompting directly simulates voting behavior using real demographic personas from the ANES dataset. V2 (§2.2.2): Time-Based Prompting enhances these personas by incorporating election-year-specific details (2016 and 2020). V3 (§2.2.3): Multi-Step Reasoning introduces an additional step to infer ideological alignment, evaluated through two methods: one using real political ideology from ANES (3rd Pipeline_Real_Ideology) and another relying on LLM-generated ideology (3rd Pipeline_Generated_Ideology).
@@
Figure 2: Comparison of the Three Pipelines on ANES 2016 and 2020. The y-axis represents the predicted voting ratio (Eq. 1). The red baseline indicates the ground truth voting ratios from the ANES dataset.
Directly prompting an LLM with persona data alone fails to accurately simulate real human voting behavior. Both the vanilla pipeline (V1) and time-based pipeline (V2) show significant distortions from the baseline, particularly favoring the Republican Party in 2016, with predicted vote shares of 63.25% (V1) and 66.38% (V2)—substantially higher than the actual proportion. Conversely, in 2020, both pipelines underestimated Democratic support, predicting 30.14% (V1) and 10.02% (V2), far below the baseline.
Introducing political ideology inference (V3) significantly improves alignment with real voting patterns. For instance, V3 predicts 46.84% Republican support in 2016 (actual: 47.7%) and 46.79% in 2020 (actual: 41.2%), demonstrating enhanced accuracy. Notably, in 2016, LLM-generated ideology in V3 slightly outperforms the original ANES ideology, suggesting LLMs can generate meaningful ideological features. [^4] Due to missing demographic variables in the 2020 ANES dataset, we conducted ideology generation only on 2016. These findings provide strong evidence that our multi-step pipeline effectively simulates human decision-making using real-world persona data.
2.3.2 2020 U.S. Election Simulation: State-Level Evaluation
Building on the successful validation of our proposed framework on the ANES dataset, we scale up the simulation with synthetic persona data designed to reflect the U.S. population distribution. Specifically, we use GPT-4o to simulate the 2020 U.S. presidential election, selecting five traditionally Republican states, five traditionally Democratic states, and 11 competitive (swing or tipping-point) states for state-level simulations. We then compare the predicted outcomes with official results from the Federal Election Commission (FEC).
To effectively measure simulation accuracy, we introduced two metrics: Weighted Absolute Error (WAE) (Eq. A4) and Weighted Mean Squared Error (WMSE) (Eq. A5). WAE measures overall alignment between simulated and actual outcomes, while WMSE assigns greater penalties to larger deviations due to its squared formulation. Together, these metrics provide a comprehensive and robust evaluation of aggregate simulation accuracy.
GPT-4o Simulation of 2020 U.S. Election
Metric V1 (%) V2 (%) V3 (%)
WAE 22.78 14.97 5.24
WMSE 5.46 2.34 0.37
@@
Table 1: Comparison of simulation accuracy metrics across three pipelines for 2020 U.S. election (GPT-4o).
As shown in Table 1, and consistent with the public benchmark evaluation, V1 and V2 exhibited significant distortions from actual outcomes due to higher WAE and WMSE values. In contrast, V3 demonstrated substantially greater accuracy, achieving 5.24% WAE and 0.37% WMSE, indicating a closer alignment with real voting behavior. At the state level, V3 correctly predicted outcomes in all traditionally Republican and Democratic states and 9 of 11 swing states, with only minor deviations in North Carolina (NC) and Arizona (AZ). A detailed breakdown of state-level results is provided in Appx. B. These findings underscore V3’s ability to model complex voter dynamics and closely reflect real-world electoral trends.
2.3.3 2024 U.S. Election Simulation: State-Level Cross-Model Evaluation
Evaluating only the 2020 U.S. election simulation results risks conflating an LLM’s ability to simulate voter behavior with its memorization of well-documented election outcomes Wang et al. (2024). Since the 2020 election is widely covered in most LLMs’ pretraining corpora, results may reflect recall rather than true generalization. To rigorously assess LLMs’ ability to generalize to unseen data—and to evaluate the robustness of our multi-step pipeline across models—we conducted extensive simulations for the 2024 U.S. election using LLMs trained on corpora predating 2024.
Our primary simulation used GPT-4o with the multi-step reasoning pipeline (V3) to predict voting outcomes across all 50 U.S. states. To enable cross-model comparisons while optimizing computational resources, we further evaluated multiple models on the 11 swing and tipping-point states analyzed in the 2020 simulation. This evaluation included three GPT-4o pipelines (V1, V2, V3), three LLaMA 3.1 70B pipelines (V1, V2, V3), and the V3 pipelines for Qwen 72B and DeepSeek-V3. Additionally, we compared these results with an existing LLM-based election prediction study Zhang et al. (2024b). The overall cross-model results are summarized in Table 2.
Multi-LLM Simulation of 2024 U.S. Election
Metric GPT-4o V1 GPT-4o V2 GPT-4o V3 LLaMA3-70B V1 LLaMA3-70B V2 LLaMA3-70B V3 Qwen-72B V3 DeepSeek-V3 V3 Zhang et al., 2024
WAE 21.35 25.96 3.49 19.97 10.23 6.88 10.66 14.57 4.70
WMSE 4.83 6.89 0.22 4.15 1.42 0.68 1.47 2.43 0.30
BM -21.35 -25.96 -2.95 -19.97 -10.23 -5.44 -9.47 -14.57 1.26
Table 2: Evaluation metrics for the 2024 U.S. election simulations across different LLMs and pipelines.
Consistent with the 2020 election simulation, our theory-driven multi-step pipeline (V3) outperformed V1 and V2 within the same LLM, enabling more accurate simulations of human voting behavior. Notably, GPT-4o achieved the lowest errors with 3.49% WAE and 0.22% WMSE, while LLaMA 3.1-70B followed with 6.88% WAE and 0.68% WMSE. These results confirm that our multi-step approach enhances LLMs’ ability to produce human-like voting simulations compared to single-prompt methods. To examine systematic differences across models, we further tested V3 on Qwen-72B and DeepSeek-V3. The cross-model evaluation showed that GPT-4o’s simulation aligned most closely with real human voting behavior, demonstrating its superior ability to capture voter dynamics. A detailed breakdown of state-level simulation results for 2024 election is provided in Appx. C.
3 Beyond Accuracy: Limitations in Large-Scale Political Simulations (RQ3)
In §2.2.3, we introduced a multi-step reasoning pipeline to enhance LLMs’ ability to simulate human voting behavior. However, human decision-making is complex and uncertain Treier and Hillygus (2009), and LLMs may struggle to fully capture its nuances. In the following section, we examine the challenges and constraints LLMs face in simulating real-world decision processes, offering insights to guide future research on LLM-based human behavior modeling.
We focus on three key issues: (1) the systematic political bias in LLMs originating from pretraining data (§3.1); (2) the reinforcement of demographic stereotypes (§3.2), and (3) the model’s tendency to overestimate the influence of certain predictors on decision-making outcomes (§3.3).
3.1 Systematic Political Bias in LLMs
Previous research has shown that LLMs exhibit varying ideological leanings due to biases in their pretraining data Feng et al. (2023). To evaluate whether these tendencies affect LLM-based human behavior simulations, we introduce a new metric: Bias Metric (BM) (Eq. A6). BM quantifies systematic bias by measuring whether simulated personas consistently favor one party over the other. Specifically, a BM $>$ 0 indicates a Republican-leaning bias, while a BM $<$ 0 suggests a Democratic-leaning bias.
As shown in Table 1 and 2, LLMs using single-prompt persona-based approaches (V1 and V2) exhibit a strong Democratic bias. Ours (V3) reduces this bias but does not fully eliminate it—lowering BM from --21.35% to --2.95% in GPT-4o and from --19.97% to --5.44% in LLaMA 3.1-70B.
Furthermore, systematic affiliations vary across models. When tested on unseen data using pipeline V3, DeepSeek-V3 displayed the strongest Democratic bias (--14.57%), while GPT-4o showed the smallest (--2.95%).
@@ 3.1 Systematic Political Bias in LLMs
Future Direction 1: Addressing Embedded Political Skewness in Pretrain Corpora. The persistent Democratic skew in simpler pipelines and the residual bias in V3 suggest deeper imbalances in the pretraining corpus. These biases may stem from uneven representation of political perspectives or disproportionate exposure to certain ideologies in the training data Jenny et al. (2024). Mitigating this issue requires a comprehensive approach, including analyzing corpus composition, adopting balanced data selection strategies, and implementing model-level interventions such as adversarial debiasing or targeted prompt engineering. Addressing these root causes will help future simulation studies produce more balanced and reliable results, strengthening LLMs as tools for political analysis and decision-making Li et al. (2024).
3.2 Reinforcement of Demographic Stereotypes
Beyond systematic bias, it is crucial to examine whether LLM simulations capture real-world demographic voting patterns. We focus on four key demographic dimensions—gender, ethnicity, age, and education—highlighted in Pew Research Center’s 2020 study, Behind Biden’s 2020 Victory Center (2021), which identified systemic voting preferences across different groups (e.g., men leaning more Republican than women, and white voters showing stronger Republican support than other ethnic groups).
To evaluate whether these demographic trends emerge in LLM simulations, we compared GPT-4o’s 2020 predictions with Pew’s 2020 findings. As shown in Figure 3, our multi-step pipeline more accurately replicates real-world demographic patterns than direct prompting. However, the LLM also amplifies these patterns, exaggerating intra-group voting tendencies. For example, among male voters, the actual Republican preference gap is 2%, but the LLM-simulated male personas exhibit a 47.6% Republican bias. Similar overamplification appears across race, age, and education categories (see Appx. D).
These findings reveal that LLMs impose systematic stereotypes, amplifying intra-group similarities and oversimplifying the complexity of real human decision-making.
Figure 3: Comparison of LLM-simulated voting patterns by gender and race against real human data from the Pew Report.
@@ 3.2 Reinforcement of Demographic Stereotypes
Future Direction 2: Mitigating Demographic Stereotypical Biases. While the LLM’s ability to capture directional trends from real-world data is promising, its overemphasis on demographic distinctions raises concerns Chang et al. (2024). Such exaggerations risk reinforcing stereotypes and misrepresenting demographic groups, potentially distorting analytical insights. Future research should focus on calibrating LLM outputs, refining prompt designs, and integrating counterbalancing information to ensure simulations are both directionally accurate and proportionally realistic Park et al. (2024). Maintaining fairness in LLM-based simulations—rather than amplifying biases—is essential for developing ethical and reliable computational social science tools.
3.3 Overestimated Influence of Political Ideology on Voting Preference
To assess the validity of our multi-step reasoning framework in aligning with real human voting behavior, we examine the extent to which political ideology predicts voting preference. Specifically, we run a logistic regression using LLM-inferred ideology (on a 1 to 7 scale, where 1 = extremely liberal and 7 = extremely conservative) as the predictor and voting preference (0 = Democrat, 1 = Republican) as the outcome. We compare the regression coefficients and pseudo R-squared values with those from a logistic regression on real human data from ANES.
Our results (Figure 4) confirm that ideology strongly correlates with voting preference, with liberals favoring Democrats and conservatives leaning Republican. However, this relationship may be exaggerated in LLM simulations, as evidenced by the higher regression coefficients ($\beta$) and $R^{2}$ values in GPT-4o simulations compared to real human data. Specifically, the regression coefficients and goodness-of-fit metrics for GPT-4o simulations (2024: $\beta=4.95,R^{2}=0.75$; 2020: $\beta=7.76,R^{2}=0.91$) exceed those observed in actual human responses from ANES ($\beta=1.53,R^{2}=0.44$). Additionally, for the 2024 election simulations using LLaMA, Qwen, and DeepSeek, we observe R-squared values approaching 1, indicating complete separation—a condition where the predictor perfectly predicts the outcome, causing the model to fail to converge. Due to this instability, we exclude these models from visualization.
Future Direction 3: Human-in-the-Loop Reinforcement Learning. This finding aligns with the concept of “hyper-accuracy distortion” Aher et al. (2023), where LLMs improve reasoning by closely following ideological patterns but risk exaggerating predictive certainty. Since human decision-making is inherently complex and dynamic Treier and Hillygus (2009), developing an effective human-in-the-loop RL framework Zhang and Lu (2024) is crucial. Such a framework would iteratively refine LLM behavior through human feedback, enabling more nuanced and realistic simulations of human decision-making. Advancing this approach presents a promising avenue for future research, bridging the gap between LLM reasoning patterns and real-world human behaviors.
Figure 4: Logistic regression analysis of political ideology and voting preference, comparing LLM simulations with real human data (ANES).
@@ 4 Related Work
4.1 LLMs in Political Science
Recent research has examined the use of large language models in debates, election forecasting Taubenfeld et al. (2024); Jiang et al. (2024), and legislative behavior Baker and Azher (2024). However, these models have also been found to exhibit inherent biases Feng et al. (2023), and some argue that political neutrality is unattainable Fisher et al. (2025). While earlier frameworks have detailed the strengths and limitations of large language models in generative and predictive tasks, their application to modeling voter behavior remains limited. Thus, our study introduces a large-scale simulation framework that incorporates social science theories to more accurately model political decision-making.
4.2 Simulating Human Decision-Making
LLM-based simulations of human behavior draw insights from public opinion, economics, and social psychology. Chuang et al. (2024) modeled opinion dynamics and polarization, whereas Ross et al. (2024) applied utility theory to capture economic decision-making patterns. Zhang et al. (2024a) examined LLMs’ capability to simulate collaboration and conformity, though Chang et al. (2024) noted that these models tend to overestimate political homophily in social networks. Our approach builds on these findings by integrating established political science theories, such as ideological sorting and partisanship, to enhance the realism of voter behavior simulations.
5 Conclusion
In this work, we introduced a theory-driven multi-step reasoning pipeline that combines demographic, time-sensitive, and ideological information to simulate voter decision-making. Our evaluations on benchmark and state-level datasets show that our approach improves prediction accuracy and reduces bias compared to simpler methods, demonstrating that large language models can replicate key aspects of human voting behavior and provide a useful tool for research in political science.
Limitations
Our approach has three main limitations. First, our time-dependent modeling does not capture dynamic factors such as changes in public opinion, shifts in media narratives, or unexpected events; incorporating these elements would require real-time data integration, which is beyond the scope of this study. Second, while our experiments include different LLMs, the ideology-based framework’s ability to generalize to more complex, multi-party scenarios—such as those in countries like Japan or France—remains untested due to time and resource constraints. Third, despite efforts to mitigate systematic biases through multi-step reasoning, LLMs still exhibit residual political skewness and exaggerate the influence of ideological alignment on voting behavior. Addressing these biases may require further human-centered refinement efforts.